OpenStudy (dan815):

how do you do this If the height of the cone is 5 and the base radius is 4, write a parameterization of the cone in terms of r=s and θ=t. x(s,t)= , y(s,t)= , and z(s,t)= , with ≤s≤ and ≤t≤ .

OpenStudy (dan815):

@wio

OpenStudy (anonymous):

 Like this?

OpenStudy (dan815):



OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

Okay so first lets do \(z\) cause it is easiest. We want to convert the radius to height.

OpenStudy (anonymous):

We have two points: \((s,z)=(0,5)\) and \((s,z)=(4,0)\)

OpenStudy (anonymous):

The equation of a line that goes through the points \((x_1,y_1)\) and \((x_2,y_2)\) is given by: \[ y-y_1 = \frac{y_2-y_1}{x_2-x_1} (x-x_1) \]

OpenStudy (anonymous):

Use this if formation to find \(z(s)\)

OpenStudy (anonymous):

\(z(s,t)=z(s)\) since \(z\) will be constant with respect to \(t\).

OpenStudy (anonymous):

We also know that \(0\le s \le 4\)

OpenStudy (anonymous):

@dan815 You following me?

OpenStudy (dan815):

ya im reading

OpenStudy (dan815):

i dont understand that equation of a linne

OpenStudy (dan815):

will i get the same answer with y = mx+b

OpenStudy (dan815):

oh nvm now i see go on

OpenStudy (anonymous):

Basically \(z(s,t)\) is a line. \(z(s,t)=ms+b\)

OpenStudy (anonymous):

So can you find \(z\) to start with?

OpenStudy (dan815):

so (5/-4)x + 5?

OpenStudy (anonymous):

yes, but not \(x\), it's \(s\)

OpenStudy (dan815):

right

OpenStudy (anonymous):

do you remember the parametrization of a circle?

OpenStudy (dan815):

x^2+y^2 = 16?

OpenStudy (dan815):

sintheta^2+cos^2

OpenStudy (dan815):

sin(t)^2+cos(t)^2=16 to sin(t)^2+cos(t)^2=0

OpenStudy (anonymous):

It's \((x(t), y(t)) = (\cos(t),\sin(t))\)

OpenStudy (dan815):

how do u write an equation where that circle is decreasing as z increases

OpenStudy (anonymous):

But this is for the unit circle, if you want it to have radius \(s\), we multiply that in: \[ x(s,t)= s\cos(t)\\ y(x,t)=s\sin(t) \]

OpenStudy (anonymous):

\[ x(s,t)= s\cos(t)\\ y(s,t)=s\sin(t) \]

OpenStudy (dan815):

why does that work

OpenStudy (anonymous):

Remember the \(z\) changes is because \(s\) that is increasing.

OpenStudy (anonymous):

It works because of the equation of a circle: \[ x^2+y^2= r^2 \\ [s\cos(t)]^2+[s\sin(t)]^2=r^2\\ s^2\cos^2(t)+s^2\sin^2(t)=r^2\\ s^2(\cos^2(t)+\sin^2(t))=r^2\\ s^2= r^2 \]

OpenStudy (dan815):

ok gotcha

OpenStudy (dan815):

thanks

OpenStudy (anonymous):

Remember the bounds of \(s\) and \(t\)

OpenStudy (dan815):

0 - 4 and 0-2pi

OpenStudy (anonymous):

good

OpenStudy (dan815):

ty teacher :D